How many more boxes would be needed to package 1200 magazines in boxes of 10 than in boxes of 12? Explain how you derived your answer.
1
Expert's answer
2019-09-06T07:32:38-0400

To package 1200 magazines in boxes of 10 we need 1200/10 = 120 boxes.

To package 1200 magazines in boxes of 12 we need 1200/12 = 100 boxes.

So, we need 120 - 100 = 20 more boxes to package 1200 magazines in boxes of 10 than in boxes of 12.